OpenStudy (anonymous):

-ln x = t/50 + C Solve for x

OpenStudy (anonymous):

\[-\ln x = \left(\begin{matrix}t \\ 50\end{matrix}\right)+ C\]

OpenStudy (anonymous):

\[ x = -e^{t/50} - e^c\]

OpenStudy (anonymous):

Is the e when you divie the - ln ?

OpenStudy (anonymous):

no, to get x out of ln, we take the e of ln x, so e and ln canceles, leaving x on left side on the right side, every term becomes exponent of the e

OpenStudy (anonymous):

Do I add the exponents and leave a postive e? Answer: \[e ^{-C-t/50}\]

OpenStudy (anonymous):

yes you can simplify like that